OpenStudy (anonymous):

Giving out medal. Solve the system using substitution. Check your solution. 2x-y=73 5y=x The solution is = ( , ) Type an ordered pair, simplify your solution, Type an integer or fraction.

OpenStudy (campbell_st):

so replace x with 5y in the 1st equation \[2 \times 5y - y = 73\] now solve for y. when you get the answer, just substitute into either equation to find x. hope it helps p.s. you might like to check its 73 and possibly not a typo for 72

OpenStudy (anonymous):

Thanks I think I was missing a step in the equation. I thought it said "solve the system using elimination". Thank you though

OpenStudy (anonymous):

@campbell_st Can you answer my pm?

OpenStudy (campbell_st):

ok... after doing the substitution you get 10y - y = 73 or 9y = 73 which is a bit odd, 72 would have been a better number

OpenStudy (anonymous):

Yea It says 73 in the problem. Yea it would've been better because 9*8=72.